Skip to end of metadata
Go to start of metadata

Sonics and gas analyzers

No known problems at this time. Unlike the RMYoungs they replaced, the METEKs at tnw07b 4m and v07 8m do not report ldiag, so those are all missing in the QC tables, but I don't know at the moment how to fix that.

TRH Sensors

The only sensor known to be down at this point is the 60m on tse11.

Other issues

v05

We visited v05 to investigate why it was offline, even though the Ubiquiti was still reachable. Eventually we rebooted the Ubiquiti and the connection came back. The DSM had been up the whole time so no data were lost.

tnw05 rsync

The tnw05u radio continues to be a problem, hanging up rsync connections. No explanation or fix yet.

v04

Tcase.in is still down and nothing has been done about it. We still need to work on mounting a NR01 on the dark horse with the KZ.

 

  • No labels